OpenStudy (anonymous):

convert the following complex number to polar (trigonometric) form. 4+3i Please don't just give me the answer, I would like to know how to get the answer

OpenStudy (anonymous):

wanna guess whats the radius

OpenStudy (anonymous):

what does it have to do with radius?

OpenStudy (anonymous):



OpenStudy (anonymous):

polar coordinates need a radius

OpenStudy (anonymous):

rsintheta, rcostheta

OpenStudy (anonymous):

imagine its plotted real vs imaginary

OpenStudy (anonymous):

I havent learn any of these. In my textbook, it tells me: z=a+bi z=r(cos # + isin #) where z = |v| = root(a^2 + b^2) and tan # = b/a

OpenStudy (anonymous):

r is the radius here

OpenStudy (anonymous):

r can be obtained by taking root of (a^2+b^)

OpenStudy (anonymous):

so it will be 5 in your case

OpenStudy (anonymous):

oh I even wrote it haha

OpenStudy (anonymous):

JOKES

OpenStudy (anonymous):

u hacked ur self =)

OpenStudy (anonymous):

but how can i find the sin# and cos#

OpenStudy (anonymous):

so getting theta should be a piece of cake too

OpenStudy (anonymous):

tan theta= b/a so its tan theta=3/4 theta=36.9 deg

OpenStudy (anonymous):

cmon man gr1 math u got this

OpenStudy (anonymous):

=P

OpenStudy (anonymous):

sorry i dont mean to be condescending do you still have a question??

OpenStudy (anonymous):

dw if u didnt get it ill be happy to explain

OpenStudy (anonymous):

(i realy wont be)

OpenStudy (anonymous):

I think i got it

OpenStudy (anonymous):

kk cool

OpenStudy (anonymous):

i even wrote it

OpenStudy (anonymous):

letme just confirm with you

OpenStudy (anonymous):

z = a + bi tan = b/a y question is 4+3i so tan = 3/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use ur calculator now and find the angle tan^-1

OpenStudy (anonymous):

arc tan 3/4 = 36.9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would the final answer be 36.9?

OpenStudy (anonymous):

r sin theta should be 4 r cos theta should be 3 so by now you should realized that teacher hacked u

OpenStudy (anonymous):

well i dont know how you wanna leave it

OpenStudy (anonymous):

the question is just "Convert the following complex numbers to pola r (trigonometric) form.

OpenStudy (anonymous):

(4,3) or 5 sin(36), 5cos(36) or 5 radius, theta=36 are all valid ways of displaying the answer

OpenStudy (anonymous):

the thing is, there are like 5 of these questions and this will be on the exam tomorrow lol so it's confusing on which answer to put tomorrow

OpenStudy (anonymous):

your teacher will accept the first one for sure if there is some discrepancy you can ask teacher before test what format she wants it in

OpenStudy (anonymous):

this is what I'm assuming she wants it in form z=r(cos# + isin#) since that's what the textbook states as the polar form

OpenStudy (anonymous):

what do u think?

OpenStudy (anonymous):

yeah but if you simplify it, it just comes down to 4,3 =P

OpenStudy (anonymous):

but yeah do what the textbook tells u to

OpenStudy (anonymous):

maybe it's just solve for z using that form?

OpenStudy (anonymous):

lol go study something else dont over think it good night get a good sleep

OpenStudy (anonymous):

nop I GOT IT! i read the next page of my textbook and it's basically just writing it as z=r(cos#+isin#).

OpenStudy (anonymous):

while filling in each of the things (z, r, #)